Okay, the instructions say to Solve the Radical Equation...
6-√2x=-2
I know how to solve Radical Equations, but we were not told how to do ones that had a negative sign before the √. I really appreciate your help.

Answer by ilana(307)

If your only problem is the negative, just remember that the square root of a number is still just another number. So anything you can do to a number, you can do to the square root of the number. By that I mean the following:
If you had the equation 6-2x=2, you could just add 2x to both sides in order to get rid of the negative sign. So that would become 6=2x+2.
In the problem you gave, 6-√2x=-2 becomes 6=-2+√2x, which you said you know how to solve. Good luck!